Newham Corporation produces and sells two products. In the most recent month, Product R10L had sales of $23,000 and variable expenses of $10,070. Product X96N had sales of $36,000 and variable expenses of $17,660. The fixed expenses of the entire company were $45,980. The break-even point for the entire company is closest to:
Answer #1

compute break even point for entrie company

Particular

product

R10L

Product

X96N

Total
Sales $23000 $36000 $59000
(-) variable expenses ($10070) ($17660) ($27730)
CONTRIBUTION $12930 $18340 $31270
CONTRIBUTION MARGIN ratio=CONTRIBUTION/sales×100 $12930/$23000×100=56% $18340/$36000×100=51% 53%
Fixed expenses $45980

  break even point for entrie company= fixed expenses/CONTRIBUTION MARGIN ratio

   =$45980/53%=$86755

   ALL THE BEST

PLEASE DO SUPPORT US

   ANY DOUBT PLEASE COMMENT BELOW

THANK YOU